OpenStudy (anonymous):

Write the simplest polynomial function with the given roots: 2i, sqrt(3), 4. I have gathered that a complex conjugate is needed for 2i, i.e. (x+2i)(x-2i) and the 4 translates into (x-4). Does the sqrt(3) need one or two binomials?

OpenStudy (lgbasallote):

if x = 2i...then x = sqrt -2 x^2 = -2 first root is (x^2 +2) then (x^2-3) and x-4

OpenStudy (lgbasallote):

so your polynomial is (x^2+2)(x^2-3)(x-4) did you get that?

OpenStudy (lgbasallote):

oh wait sorry... x= sqrt -4 so it's (x^2 +4)

OpenStudy (accessdenied):

You also have to use the conjugate of the sqrt(3). (I think it's the irrational root theorem?)

OpenStudy (anonymous):

Yes, thank you!.